OpenStudy (anonymous):

If you were to use the substitution method to solve the following system, choose the new system of equations that would result if x was isolated in the first equation x + 6y + z = 2 x – 3y + 2z = –1 x + 5y + 3z = –4

OpenStudy (anonymous):

x = 2 - 6y - z (isolate x) now sub 2 - 6y - z in for x into the second equation x - 3y + 2z = -1 2 - 6y - z - 3y + 2z = -1 combine like terms -9y + z = -3 now sub 2 - 6y - z in for x in your third equation x + 5y + 3z = -4 2 - 6y - z + 5y + 3z = -4 combine like terms -y + 2z = -6

OpenStudy (anonymous):

thank you so much :D

OpenStudy (anonymous):

no problem :)